OpenStudy (anonymous):

expanded if the form of 2.753 is 2x1+7 x(1/10)+5x(1/100)+3x(1/1000) then what is the form of 6.324

OpenStudy (aum):

\[6.324 = 6 \times 1 + 3 \times \frac{1}{10} + 2 \times \frac{1}{100} + 4 \times \frac{1}{1000} \]